# Nonuniqueness of Leray–Hopf solutions to the unforced incompressible 3D Navier–Stokes Equation

### v1

@@ 2.2 Nonuniqueness of Leray–Hopf solutions
$\widetilde{u}(x,t)=\frac{1}{\sqrt{t}}\widetilde{U}(\frac{x}{\sqrt{t}}),\quad u^{\text{cut}}(x,t)=\frac{1}{\sqrt{t}}U^{\text{cut}}(\frac{x}{\sqrt{t}},\log t),\quad u^{\text{cor}}(x,t)=\frac{1}{\sqrt{t}}U^{\text{cor}}(\frac{x}{\sqrt{t}},\log t).$
@@
and we need to show the existence of fixed points $\mathcal{T}(U)_{s}=P_{s}U$, $\mathcal{T}(U)_{j}=P_{j}U$, for appropriately chosen $U_{j0}=P_{j}U(\cdot,0)\in X_{j}$. We remark that by Theorem 2, we have at least one unstable $j$.
@@
With the growth bound (2.13), we obtain the smoothing estimate as in Lemma 4.4 of [1]: for any $\delta>0$, $\sigma_{2}\geqslant\sigma_{1}\geqslant 0$, there exists a constant only depending on $\delta,\sigma_{2},\sigma_{1}$, such that for any $f_{s}\in X_{s}\cap H^{\sigma_{1}}$ and $\tau>0$, we have
@@
Every finite-dimensional unstable eigenspace $X_{j}$ is the kernel of $(A_{j}-\lambda_{j})^{k_{j}}$ for some finite $k_{j}$. Thus it is spanned by bases $\{f_{j,k}\}$, where every $f_{j,k}$ satisfies
@@
Similarly, we have a smoothing estimate for $A_{j}$, so the bases $f_{j,k}$ are smooth.
@@
Thus $\mathcal{T}$ is a strict contraction on $\mathcal{D}_{R,\eta,U_{j0}}$. Hence, we conclude a solution to (2.11) and thus (2.10) for any small enough admissible $U_{j0}$, which contains infinitely many distinct solutions $U^{\text{cor}}$, different by their projections at time $\tau=0$, i.e. $t=1$.
@@
Finally, $\|u^{\text{cor}}\|_{L^{2}}=t^{1/4}\|U^{\text{cor}}\|_{L^{2}}\to 0$ as $t\to 0$, and we conclude the constructed $u=\widetilde{u}+u^{\text{cut}}+u^{\text{cor}}$ are infinitely many Leray–Hopf solutions with the same $L^{2}$ initial data $u_{\operatorname{loc}}$. Their smoothness implies that they are moreover suitable and (1.2) holds with equality, as in [1].
@@ 3.1.1 Simple case: $\overline{w}\geqslant 0$
First, we consider a simple case where $\overline{w}\geqslant 0$ and $L$ is thus coercive. Specifically,
@@ 3.2.3 Main theorem
The boundedness is trivial by (3.19), (3.7), and Assumption 2. By Lax-Milgram, we define the mapping $T$ by setting $T(w_{1})=w_{1}^{\ast}$, where $w_{1}^{\ast}$ solves:
@@
If $w_{1}^{\ast}\neq\widehat{w}_{1}^{\ast}$, cancel the common factor to obtain
@@ 4 Proof of Theorem 2: Analysis Part
We remark that the key challenge lies in the linear estimate (3.20), while the nonlinear and residual estimates are more straightforward. The crucial changes are that the linear operator is no longer bounded due to the term $x\cdot\nabla$, and that $H^{1}$ is not compactly embedded into $L^{2}$ on the whole space. We truncate the domain so that the associated bilinear form will be bounded, as in Assumption 2 to apply Lax-Milgram. Moreover, thanks to the gain of regularity, we show that the fixed point map maps a bounded set in $H^{1}$ to itself and is continuous in $L^{4}$. Therefore, in the truncated domain, we can apply the Schauder fixed-point theorem to obtain a fixed point, and taking weak limits yields the desired solution.
We first discuss the concrete setting and the linear operators in Section 4.1. The finite-rank approximation for a general type of compact perturbations applicable to $U$ and $v$ simultaneously is discussed in Section 4.2. Proposition 1 is addressed in Section 4.3, as Proposition 2 follows a similar verification process in Section 4.4. Section 4.5 briefly presents our numerical results, providing evidence for the existence of the solution. Implementation details of the numerical algorithm are deferred to Sections 5 and 6.
@@ 4.1 Setting and decomposition of linear operators
$Q_{s}(x)+\eta_{1}I_{3}=Q_{p}(x)-Q(x),\quad Q_{p}(x)\geqslant 0,\quad Q(x)\geqslant 0.$
@@
Therefore, the negative part $Q(x)$ of $Q_{s}(x)$ has compact support. Using $\nabla\overline{U}=Q_{a}+Q_{p}-\eta_{1}I_{3}-Q$, the operator $L_{U}$ is finally decomposed as:
$L_{1}^{U}U$ $:=(-\eta_{1}-\frac{1}{2})U-\frac{1}{2}x\cdot\nabla U-\Delta U+\Pi(\overline{U}\cdot\nabla U+U\cdot(Q_{a}+Q_{p})),$ (4.1)
$L_{2}^{U}U$ $:=-\Pi(U\cdot Q).$ (4.2)
With this refined decomposition, we explicitly verify coercivity:
$\langle L_{1}^{U}U,U\rangle=(\frac{1}{4}-\eta_{1})\|U\|_{L^{2}}^{2}+\|\nabla U\|_{L^{2}}^{2}+\langle Q_{p}U,U\rangle\geqslant(\frac{1}{4}-\eta_{1})\|U\|_{L^{2}}^{2}+\|\nabla U\|_{L^{2}}^{2}.$ (4.3)
@@
Here $\lambda^{\text{lin}}$ is linear in $v$, while $\lambda^{\text{con}}$ is independent of $v$ with the trivial bound
$|\lambda^{\text{con}}|\leqslant\|\nabla U\|_{L^{2}}\|\overline{v}\|_{\infty}.$ (4.4)
Consequently, equation (1.17) can be decomposed as:
@@
$L^{v}v$ $:=-(\frac{1}{2}+\overline{\lambda}+\lambda^{\text{con}})v-\frac{1}{2}x\cdot\nabla v-\Delta v+\Pi(\widetilde{U}\cdot\nabla v+v\cdot\nabla\widetilde{U})-\lambda^{\text{lin}}(v)\overline{v},$
$N^{v}(v)$ $:=-\lambda^{\text{lin}}(v)v,\quad\widehat{\mathrm{E}}^{v}:=\Pi\mathrm{E}^{v}+\Pi(U\cdot\nabla\overline{v}+\overline{v}\cdot\nabla U)-\lambda^{\text{con}}\overline{v}.$
$L^{v}$ resembles $L^{U}$, except for a few additional terms in the linear operator. The $(\overline{\lambda}+\lambda^{\text{con}})v$ term provides enhanced coercivity, while the extra terms $U\cdot\nabla v$ and $v\cdot\nabla U$ are treated as perturbations. The term $\lambda^{\text{lin}}(v)\overline{v}$ should be interpreted as a projection. More explicitly, we define:
@@
Recall the definitions of $Q_{a},Q_{p},Q$ from the decomposition of $L^{U}$. Similarly, we split the linear operator $L^{v}$ into two components:
$L_{1}^{v}v$ $:=(-\eta_{1}-\frac{1}{2}-\overline{\lambda}-\lambda^{\text{con}})v-\frac{1}{2}P_{\overline{v}}(x\cdot\nabla v)-P_{\overline{v}}\Delta v+P_{\overline{v}}\Pi(\widetilde{U}\cdot\nabla v+v\cdot\nabla U+v\cdot(Q_{a}+Q_{p})),$
@@
Similarly, we will impose symmetries later for numerical purposes. We verify the coercivity of the operator $L_{1}^{v}$. Indeed, it holds that
@@
$|\langle v\cdot\nabla U,v\rangle|\leqslant\|\nabla U\|_{L^{2}}\|v\|_{L^{4}}^{2}\leqslant 2K_{4}\|\nabla U\|_{L^{2}}(\|v\|_{L^{2}}^{2}+\|\nabla v\|_{L^{2}}^{2}),$
@@
ensuring coercivity for sufficiently small $\|\nabla U\|_{L^{2}}$ and $|\lambda^{\text{con}}|$. Finally, since $Q$ is bounded of compact support, $L_{2}^{v}$ is a compact operator from $\mathcal{H}^{1}$ to $\mathcal{H}^{0}$ and bounded from $\mathcal{H}^{0}$ to $\mathcal{H}^{0}$. We have verified Assumption 2: the coercivity of $L_{U}^{1}$ and the compactness of $L_{U}^{2}$.
Finally, we establish the following bound for $\lambda^{\text{lin}}$ via an integration by parts:
$|\lambda^{\text{lin}}(v)|=|\langle\frac{1}{2}x\cdot\nabla\overline{v}-\Delta\overline{v}-\widetilde{U}\cdot\nabla\overline{v}+\overline{v}\cdot\nabla^{T}\widetilde{U},v\rangle|\leqslant M_{0}\|v\|_{L^{2}},$ (4.8)
$M_{0}:=\|P_{\overline{v}}(\frac{1}{2}x\cdot\nabla\overline{v}-\Delta\overline{v}-\overline{U}\cdot\nabla\overline{v}+\overline{v}\cdot\nabla^{T}\overline{U})\|_{L^{2}}+\|U\|_{L^{2}}\|\nabla\overline{v}\|_{\infty}+\|\nabla U\|_{L^{2}}\|\overline{v}\|_{\infty},$
where we use the orthogonality of $v$ to $\overline{v}$ and the definition of $P_{\overline{v}}$.
@@ 4.2 Finite rank approximation
Recall that we want the finite-rank operator to have a number of bases as small as possible so that numerical verification of invertibility on the space can be easier. Ideally, we should choose the eigenspace corresponding to the largest eigenvalues of $L_{2}$ so that the residue of the finite-rank approximation is small. This is not tractable in practice, however, since the eigenvectors are not explicit. We instead solve the eigenvalue problem on an explicitly constructed finite-dimensional subspace numerically, and require that $L_{2}$ acting outside of the space is small. To be precise, the construction involves three main steps:
@@
    Choose a domain $\Omega$ sufficiently large to encompass the support of $Q$, and construct a finite-dimensional collection of spectral bases on $\Omega$.
@@ 4.2.1 Construction of the finite rank approximation
We first identify a suitable domain $\Omega$ that fully encompasses the support of the matrix function $Q(x)$. Now, we introduce spectral approximations for functions defined on $\Omega$. Let $V_{N}$ denote the finite-dimensional space with the first $N$ spectral bases. We remark that the space and thus the eigenpairs associated with $U$ and $v$ will be different due to their respective parity symmetries. Consider the following $E$-inner product and its induced norm
@@
where we choose the constant $\eta_{2}$ as $0.005$. Notice that by construction of the spectral basis, the $L^{2}$-projection $I_{N}$ defined by (6.5) in Section 6 onto the space $V_{N}$ is also an $H^{1}$-projection, hence an $E$-projection.
To obtain an efficient finite-rank approximation, we consider the following eigenvalue problem within the space of spectral bases $V_{N}$:
@@ 4.2.2 Estimates of the finite rank approximation
measuring how close the approximate eigenpairs are to being exactly orthonormal in the $Q$-geometry.
We now present a lemma on quasi-orthogonality in $E$-inner product; see its proof in Appendix B.
@@
which should approximately be the eigenvalue smaller than $\overline{\lambda}_{m}$. With the bounded estimates in hand, we are in a position to present the following lemma on a rigorous estimate of the approximation error associated with $L_{\operatorname{ap}}$.

### v2

@@ 2.2 Nonuniqueness of Leray–Hopf solutions
$\widetilde{u}(x,t)=\frac{1}{\sqrt{t}}\widetilde{U}(\frac{x}{\sqrt{t}}),\quad u^{\text{cut}}(x,t)=\frac{1}{\sqrt{t}}U^{\text{cut}}(\frac{x}{\sqrt{t}},\log t),\quad u^{\text{cor}}(x,t)=\frac{1}{\sqrt{t}}U^{\text{cor}}(\frac{x}{\sqrt{t}},\log t).$ (2.7)
@@
and we need to show the existence of fixed points $\mathcal{T}(U)_{s}=P_{s}U$, $\mathcal{T}(U)_{j}=P_{j}U$, for appropriately chosen $U_{j0}=P_{j}U(\cdot,0)\in X_{j}$. We remark that by Theorem 2, we have at least one unstable eigenmode with eigenvalue $\lambda_{j}$. The Riesz projections are moreover bounded since they are of finite rank.
@@
With the growth bound (2.14), we obtain the smoothing estimate as in Lemma 4.4 of [1]: for any $\delta>0$, $\sigma_{2}\geqslant\sigma_{1}\geqslant 0$, there exists a constant depending on $\delta,\sigma_{2},\sigma_{1}$ only, such that for any $f_{s}\in X_{s}\cap H^{\sigma_{1}}$ and $\tau>0$, we have
@@
Every finite-dimensional unstable eigenspace $X_{j}$ is the kernel of $(A_{j}-\lambda_{j})^{k_{j}}$ for some finite $k_{j}$. Thus it is spanned by basis functions $\{f_{j,k}\}$, where every $f_{j,k}$ satisfies
@@
Similarly, we have a smoothing estimate for $A_{j}$, so the basis functions $f_{j,k}$ are smooth.
@@
Thus $\mathcal{T}$ is a strict contraction on $\mathcal{D}_{R,\eta,U_{j0}}$. Hence, we conclude, via the Banach fixed point theorem, a solution to (2.12) and thus (2.11) for any small enough admissible $U_{j0}$, which contains infinitely many distinct solutions $U^{\text{cor}}$, different by their projections at time $\tau=0$, i.e. $t=1$.
@@
Finally, $\|u^{\text{cor}}\|_{L^{2}}=t^{1/4}\|U^{\text{cor}}\|_{L^{2}}\to 0$ as $t\to 0$ so that $u\in L^{\infty}L^{2}$. Similarly, $\|\nabla u\|_{L^{2}}=t^{-1/4}\|\nabla U\|_{L^{2}}$ so that the $H^{1}$-norm of $u$ is bounded by $t^{-1/4}$. Thus $u\in L^{2}H^{1}$, and we conclude the constructed $u=\widetilde{u}+u^{\text{cut}}+u^{\text{cor}}$ are infinitely many Leray–Hopf solutions with the same $L^{2}$ initial data $u_{\operatorname{loc}}$. Similar to [1], we can establish their smoothness for positive times via the Duhamel representation. To be precise, we can bootstrap from $\sup_{\tau<0}e^{-\delta\tau}\|U^{\text{per}}(\cdot,\tau)\|_{H^{N}}<\infty$ to $\sup_{\tau<0}e^{-\delta\tau}\|U^{\text{per}}(\cdot,\tau)\|_{H^{N+1/2}}<\infty$ via the smoothing estimates (2.15), (2.16), and (2.13), similar to the estimate in (2.17). Their smoothness implies that they are moreover suitable and (1.2) holds with equality, as in [1].
@@
Remark 3
Our choice of localization is propagated in time via the heat smoothing. It would be interesting to explore an alternative approach based on a static spatial cutoff and direct estimates of the boundary terms generated by the cutoff. In particular, one may hope to combine such an approach with inner–outer gluing ideas used by Albritton–Brue–Colombo [3], to investigate whether an analogue of Theorem 1 could hold on smooth bounded domains (or the torus), e.g. under no-slip boundary conditions, at comparable levels of regularity. A natural starting point, similar to [3], would be to consider an ansatz of the form
$u=\eta(\tilde{u}+u^{\mathrm{cor}})+\psi,$
where $\eta$ is a cutoff, $\psi$ corrects boundary/cutoff errors, and $u^{\mathrm{cor}}$ is intended to capture the nonuniqueness mechanism associated with unstable directions. Again, it is essential to run backward for the unstable directions, so that no assumption on the smallness of positive eigenvalues is needed for the cutoff terms. Another difficulty is that the correction fields $u^{\mathrm{cor}}$ and $\psi$ are expected to satisfy a coupled system whose analysis appears more delicate in bounded geometries. We do not pursue these questions here and leave them for future work.
@@ 3.1.1 Simple case: $\overline{w}\geqslant 0$
First, we consider a simple case where $\overline{w}\geqslant 0$ and $L$ is thus coercive. Specifically, we have
@@ 3.2.3 Main theorem
The boundedness is trivial by (3.19), (3.7), and Assumption 2. By the Lax-Milgram theorem, we define the mapping $T$ by setting $T(w_{1})=w_{1}^{\ast}$, where $w_{1}^{\ast}$ solves:
@@
If $w_{1}^{\ast}\neq\widehat{w}_{1}^{\ast}$, we can cancel the common factor to obtain
@@ 4 Proof of Theorem 2: Analysis Part
We remark that the key challenge lies in the linear estimate (3.20), while the nonlinear and residual estimates are more straightforward. The crucial changes are that the linear operator is no longer bounded due to the term $x\cdot\nabla$, and that $H^{1}$ is not compactly embedded into $L^{2}$ on the whole space. We truncate the domain so that the associated bilinear form will be bounded, as in Assumption 2 to apply the Lax-Milgram theorem. Moreover, thanks to the gain of regularity, we show that the fixed point map maps a bounded set in $H^{1}$ to itself and is continuous in $L^{4}$. Therefore, in the truncated domain, we can apply the Schauder fixed-point theorem to obtain a fixed point, and taking weak limits yields the desired solution.
We first discuss the concrete setting and the linear operators in Section 4.1. The finite-rank approximation for a general type of compact perturbations applicable to $U$ and $v$ simultaneously is discussed in Section 4.2. Proposition 1 is addressed in Section 4.3, as Proposition 2 follows a similar verification process in Section 4.4. Section 4.5 briefly presents our numerical results and records the verified numerical constants used in the proof. Implementation details of the numerical algorithm are deferred to Sections 5 and 6, and a roadmap for rigorous computer-assisted proof is in Section 7.
@@ 4.1 Setting and decomposition of linear operators
$Q_{s}(x)+\eta_{1}I_{3}=Q_{+}(x)-Q_{-}(x),\quad Q_{+}(x)\geqslant 0,\quad Q_{-}(x)\geqslant 0.$
@@
Therefore, the negative part $Q_{-}(x)$ of $Q_{s}(x)+\eta_{1}I_{3}$ has compact support. Using $\nabla\overline{U}=Q_{a}+Q_{+}-\eta_{1}I_{3}-Q_{-}$, the operator $L_{U}$ is finally decomposed as:
$L_{1}^{U}U$ $:=(-\eta_{1}-\frac{1}{2})U-\frac{1}{2}x\cdot\nabla U-\Delta U+\Pi(\overline{U}\cdot\nabla U+U\cdot(Q_{a}+Q_{+})),$ (4.1)
$L_{2}^{U}U$ $:=-\Pi(U\cdot Q),$ (4.2)
where we choose $Q=Q_{-}$ to ensure the following coercivity:
$\langle L_{1}^{U}U,U\rangle=(\frac{1}{4}-\eta_{1})\|U\|_{L^{2}}^{2}+\|\nabla U\|_{L^{2}}^{2}+\langle Q_{+}U,U\rangle\geqslant(\frac{1}{4}-\eta_{1})\|U\|_{L^{2}}^{2}+\|\nabla U\|_{L^{2}}^{2}.$ (4.3)
@@
Here $\lambda^{\text{lin}}$ is linear in $v$, while $\lambda^{\text{con}}$ is independent of $v$. Consequently, equation (1.17) can be decomposed as:
@@
$L^{v}v$ $=-(\frac{1}{2}+\overline{\lambda}+\lambda^{\text{con}})v-\frac{1}{2}x\cdot\nabla v-\Delta v+\Pi(\widetilde{U}\cdot\nabla v+v\cdot\nabla\widetilde{U})-\lambda^{\text{lin}}(v)\overline{v},$ (4.4)
$N^{v}(v)$ $=-\lambda^{\text{lin}}(v)v,\quad\widehat{\mathrm{E}}^{v}=\Pi\mathrm{E}^{v}+\Pi(U\cdot\nabla\overline{v}+\overline{v}\cdot\nabla U)-\lambda^{\text{con}}\overline{v}.$
We observe that $L^{v}$ resembles $L^{U}$, except for a few additional terms in the linear operator. The $(\overline{\lambda}+\lambda^{\text{con}})v$ term provides enhanced coercivity, while the extra terms $U\cdot\nabla v$ and $v\cdot\nabla U$ are treated as perturbations. The term $\lambda^{\text{lin}}(v)\overline{v}$ can be interpreted as a projection. More explicitly, we define:
@@
Recall the definitions of $Q_{a},Q_{+},Q$ from the decomposition of $L^{U}$. Similarly, we split the linear operator $L^{v}$ into two components:
$L_{1}^{v}v$ $:=(-\eta_{1}-\frac{1}{2}-\overline{\lambda}-\lambda^{\text{con}})v-\frac{1}{2}P_{\overline{v}}(x\cdot\nabla v)-P_{\overline{v}}\Delta v+P_{\overline{v}}\Pi(\widetilde{U}\cdot\nabla v+v\cdot\nabla U+v\cdot(Q_{a}+Q_{+})),$
@@
Similarly, we will impose symmetries later for numerical purposes. We verify the coercivity of the operator $L_{1}^{v}$. Indeed, recalling that $\overline{\lambda}$ is negative and $|\lambda^{\text{con}}|$ is small, it holds that
@@
$|\langle v\cdot\nabla U,v\rangle|\leqslant\|\nabla U\|_{L^{2}}\|v\|_{L^{4}}^{2}\leqslant 2K_{4}\|\nabla U\|_{L^{2}}(\|v\|_{L^{2}}^{2}+\|\nabla v\|_{L^{2}}^{2}),$ (4.7)
@@
ensuring coercivity for sufficiently small $\|\nabla U\|_{L^{2}}$ and $|\lambda^{\text{con}}|$. Finally, since $Q$ is bounded of compact support, $L_{2}^{v}$ is a compact operator from $\mathcal{H}^{1}$ to $\mathcal{H}^{0}$ and bounded from $\mathcal{H}^{0}$ to $\mathcal{H}^{0}$. We have verified Assumption 2: the coercivity of $L_{v}^{1}$ and the compactness of $L_{v}^{2}$.
@@ 4.2 Finite rank approximation
Recall that we want the finite-rank operator to have a number of basis functions as small as possible so that our numerical verification of invertibility on the space can be easier. Ideally, we should choose the eigenspace corresponding to the largest eigenvalues of $L_{2}$ so that the residue of the finite-rank approximation is small. This is not tractable in practice, however, since the eigenvectors are not explicit. We instead solve the eigenvalue problem on an explicitly constructed finite-dimensional subspace numerically, and require that $L_{2}$ acting outside of the space is small. To be precise, the construction involves three main steps:
@@
    Choose a domain $\Omega$ sufficiently large to encompass the support of $Q$, and construct a finite-dimensional collection of spectral basis functions on $\Omega$.
@@ 4.2.1 Construction of the finite rank approximation
We first identify a suitable domain $\Omega$ that fully encompasses the support of the matrix function $Q(x)$. Now, we introduce spectral approximations for functions defined on $\Omega$. Let $V_{N}$ denote the finite-dimensional space with the first $N$ spectral basis functions. We remark that the space and thus the eigenpairs associated with $U$ and $v$ will be different due to their respective parity symmetries. Consider the following $E$-inner product and its induced norm
@@
where we choose the constant $\eta_{2}$ to be $0.005$. Notice that by construction of the spectral basis, the $L^{2}$-projection $I_{N}$ defined by (6.5) in Section 6 onto the space $V_{N}$ is also an $H^{1}$-projection, hence an $E$-projection.
To obtain an efficient finite-rank approximation, we consider the following eigenvalue problem within the space of spectral basis functions $V_{N}$:
@@ 4.2.2 Estimates of the finite rank approximation
measuring how close the approximate eigenpairs are to satisfying exact orthonormality in the $Q$-geometry.
We now present a lemma on the quasi-orthogonality in the $E$-inner product; see its proof in Appendix B.
@@
which should approximately be the eigenvalue smaller than $\overline{\lambda}_{m}$. In practice, since we only need to provide an upper bound of $K_{6}$, we can increase $K_{6}$ a bit and only need to ensure and verify a positive-definiteness in the following sense:
$\sup_{u\in V_{N}}K_{6}\|u\|^{2}_{E}+\sum_{j=1}^{m}\overline{\lambda}_{j}^{-1}\langle u,Q\overline{\psi}_{j}\rangle_{\Omega}^{2}-\langle u,Qu\rangle_{\Omega}\geq 0,$ (4.17)
which we can verify in the matrix representation via a numerical $LDL^{T}$ decomposition and by showing the residue is small and diagonal entries are away from 0.
With the bounded estimates in hand, we are in a position to present the following lemma on a rigorous estimate of the approximation error associated with $L_{\operatorname{ap}}$.